# Multimodal Contrastive Representation Learning in Augmented Biomedical Knowledge Graphs

### v1

@@ 4.4 Graph Contrastive Learning
4.5 Link Prediction in KG Embedding
KG Embedding involves an embedding function $e:E\cup R\rightarrow\mathcal{X}$, which maps entities and relations in a knowledge graph to elements within an embedding space $\mathcal{X}$. Additionally, it includes a scoring function $f:\mathcal{X}^{3}\rightarrow\mathbb{R}$ that, given the embeddings of entities and relations in a triple, computes a score indicating the likelihood or validity of the triple. In our experiment, we utilize Relational Graph Convolutional Network (RGCN) Schlichtkrull et al. (2017) as the encoder to extract embeddings from graph-structured data that includes relational information. We then employ DistMult Yang et al. (2015) as a scoring function to map entities and relations to vector scores. The essence of the link prediction task lies in classifying the existence of edges between entities, where positive edges are drawn from the dataset and negative edges are randomly sampled. Binary Cross Entropy (BCE) loss is employed to evaluate the effectiveness of the classification as follows:
@@ 4.5 Link Prediction in KG Embedding
$\mathcal{L}_{\text{reg}}=\lambda\left(\|\mathbf{X}\|^{2}+\|\mathbf{R}\|^{2}\right),$
@@
To facilitate effective batch-wise training, we utilize the GraphSAINT sampling method Zeng et al. (2020) for graph sampling. This approach employs the Random Walks technique to sample subgraphs, ensuring the presence of existing edges within each batch for the link prediction task.
5 Experiments
5.1 Experimental Setup
@@
In our experiments, we utilize two principal datasets: PrimeKG++ and the DrugBank drug-target interaction dataset Knox et al. (2024a). PrimeKG++ serves as our primary dataset, enriched with detailed attribute information across a variety of biological entities, making it highly suitable for comprehensive model training and evaluation. The DrugBank dataset, a curated biomedical knowledge graph, focuses specifically on drug-target protein interactions. It comprises 9,716 FDA-approved drugs and 846 protein targets, encompassing a different set of relations and nodes compared to PrimeKG++. However, the DrugBank dataset originally lacked node attributes, necessitating augmentation by incorporating detailed attribute information akin to that in PrimeKG++, thereby ensuring a comprehensive evaluation and robust performance of our model. By leveraging the enriched attribute information integrated into both datasets, we aim to thoroughly evaluate our framework’s ability to handle both broad and domain-specific biomedical knowledge graphs, enabling a rigorous assessment of its performance and generalizability.
@@
With the introduction of PrimeKG++, our augmented dataset, we conducted a comprehensive evaluation of our approach by exploring a variety of widely-used configurations. We experimented with three well-established Graph Contrastive Learning (GCL) models: Graph Gaussian Diffusion (GGD), Graph Random Coloring Ensemble (GRACE), and Deep Graph Infomax (DGI). Additionally, we examined different attribute fusion methods, including Attention Fusion and Relation-guided Dual Adaptive Fusion (ReDAF), which weigh each modality differently before fusion. As a baseline, we also included a simple fusion approach ("None") where embeddings from various modalities were combined using a mean operation without explicit weighting. To provide additional context, we compared these configurations against models trained with Random Initialization and direct Language Model (LM)-derived embeddings. Rather than focusing on identifying a single optimal configuration, our objective was to demonstrate the versatility and robustness of the proposed approach across widely-used methods. We experimented with different configurations to showcase how our framework can be applied in diverse settings. While the choice of components may depend on the specific characteristics of the dataset, our intention was to highlight the adaptability of our framework, ensuring it performs effectively under multiple configurations.
@@
The reported results are based on models with the lowest validation loss observed during training, evaluated over 100 epochs. The statistics of the dataset splits are summarized in Table 1. Our model implementations are built using PyTorch and trained on a single NVIDIA A100 GPU over 3 hours for training. Detailed settings for all hyperparameters and summary of our models are provided in Table 4 and Table 6 respectively in Table 4. This setup ensures a rigorous and reproducible evaluation framework for assessing the performance and generalizability of our proposed methods.
Table 1: Statistics of Triple Splits for PrimeKG++ and DPI Benchmark.
@@
5.2 Results and Discussion
5.2.1 PrimeKG++
@@
Further empirical studies on edge-wise precision under more robust conditions, detailed in Section E.2, underscore our framework’s ability to generalize effectively across diverse scenarios.
5.2.2 DrugBank DTI
@@
Table 2: Link prediction performance on the PrimeKG++ dataset with varying negative sampling ratios.
Initial Embedding GCL Models Attribute Fusion 1:1 1:3 1:5
@@
Direct LM-derived - None 0.993 0.975 0.982 0.934 0.972 0.902
@@
Table 3: Link prediction performance on the DrugBank DTI dataset with varying negative sampling ratios.
Initial Embedding GCL Models Attribute Fusion 1:1 1:3 1:5
@@
Direct LM-derived - None 0.994 0.957 0.988 0.884 0.982 0.822
@@
5.3 Latent Space Visualization of Embeddings
@@
Figure 3: t-SNE visualization of drug embeddings for a single protein with the highest number of interactions in the PrimeKG++ dataset. The left panel displays embeddings derived solely from the Language Model (LM), while the right panel shows embeddings generated using our proposed approach (GRACE + ReDAF). Drugs interacting with the selected protein are labeled in red, and non-interacting drugs are labeled in blue. This comparison illustrates the structural differences in the latent space resulting from the two embedding methods.
The visualization results in Figure 3 reveal notable differences between the two configurations. Embeddings generated solely with Language Models (LM) showed less distinct clustering, with considerable overlap between the two groups. This overlap suggests a limited capacity to distinguish drugs interacting with the selected protein from those that do not.
In contrast, embeddings produced using our proposed method, which integrates LM with Graph Contrastive Learning (GCL), exhibited tighter clustering and more pronounced separation. This demonstrates the method’s superior ability to capture shared properties among drugs interacting with the same protein.
These results underscore the robustness of our framework in generating high-quality, interpretable embeddings that accurately represent underlying biological relationships, even when applied to unseen datasets such as DrugBank. Additional experiments on embedding quality are detailed in Section E.3.
6 Conclusion
@@ Appendix A Parameters and notations
Table 4: List of Hyperparameters
@@
Table 5: List of notations
@@
Table 6: Model summary
@@ Appendix B PrimeKG++ Creation
To obtain and integrate the metadata, in addition to PrimeKG Chandak et al. (2023), we perform the following steps:
@@
We utilize the open-source web-scraping library Selenium to fetch detailed information from authoritative sources such as Entrez Gene Maglott et al. (2010) and DrugBank Knox et al. (2024b) for each node type. This process involves the following steps: For genes/proteins, we collect both protein-coding genes, annotated with amino acid sequences, and non-coding genes, represented with nucleotide sequences. This information is fetched using the Entrez Gene API, which allows us to gather comprehensive gene data efficiently. For drugs, we source data from DrugBank, including molecular data represented as SMILES strings and antibody data identified by amino acid sequences. This comprehensive data collection ensures that we cover various aspects of drug representation. Each of these new data points is categorized into sub-types, collectively representing a node type. This detailed categorization enhances the granularity and richness of our knowledge graph.
@@ C.1 Attention Fusion
C.2 Relation-guided Dual Adaptive Fusion (ReDAF)
Given the sparse nature of PrimeKG++, we utilize the Relation-guided Dual Adaptive Zhang et al. (2024) Fusion model which produces a joint embedding projected from weighted parameters collected from individual modal training data. Additionally, the missing values of any element are consolidated with a random vector within the same vector space.
@@ Appendix D Contrastive learning settings
D.1 Deep Graph Infomax model
Deep Graph Infomax (DGI) Veličković et al. (2018) utilizes an unsupervised learning strategy for graph data by maximizing the mutual information between node representations and a global summary of the graph. The method begins with an assumption of a set of node features $X=\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N}\}$, where $N$ is the number of nodes, and each $\mathbf{x}_{i}\in\mathbb{R}^{F}$ denotes the features of node $i$. These are complemented by an adjacency matrix $A\in\mathbb{R}^{N\times N}$, which encodes the relational structure between nodes.
The core of DGI is an encoder function $E:\mathbb{R}^{N\times F}\times\mathbb{R}^{N\times N}\rightarrow\mathbb{R}^{N\times F_{0}}$ that transforms the node features and the adjacency matrix into high-level node embeddings $\{\mathbf{h}_{1},\mathbf{h}_{2},\ldots,\mathbf{h}_{N}\}$. These embeddings, or patch representations, are meant to encapsulate not only the properties of individual nodes but also their neighborhood structures.
@@ D.1 Deep Graph Infomax model
DGI employs a discriminator $D:\mathbb{R}^{F}\times\mathbb{R}^{F}\rightarrow\mathbb{R}$, which evaluates the mutual information between local patch representations and the global summary by assigning probability scores. These scores indicate how well the local patches (node embeddings) and the global summary correspond to each other in terms of information content.
For training, negative samples are generated through a stochastic corruption function $C:\mathbb{R}^{N\times F}\times\mathbb{R}^{N\times N}\rightarrow\mathbb{R}^{M\times F}\times\mathbb{R}^{M\times M}$, creating perturbed versions of the graph $(X_{e},A_{e})=C(X,A)$. The learning objective is to discriminate between the "true" patch-summary pairs and those generated from corrupted inputs using a noise-contrastive estimation with a binary cross-entropy loss.
@@
D.2 Graph Group Discrimination model
We experiment with a Group-discrimination-based method called Graph Group Discrimination (GGD) Zheng et al. (2022). Contrastive learning in this method is formulated to discriminate between groups of node embeddings, rather than individual pairs. This method leverages a binary cross-entropy loss to effectively distinguish between node samples from ‘positive’ (unaltered) and ‘negative’ (altered) graph structures.
@@ D.2 Graph Group Discrimination model
D.3 Graph Contrastive representation learning model
GRACE (Graph Contrastive Representation Learning) Zhu et al. (2020) applies stochastic augmentations to both node features and graph structure to learn robust node embeddings. For a graph with feature matrix $\mathbf{X}$ and adjacency matrix $\mathbf{A}$, two corrupted views $\mathbf{X}_{1},\mathbf{A}_{1}$ and $\mathbf{X}_{2},\mathbf{A}_{2}$ are generated by independently dropping features and edges. Node embeddings for these views are computed as $\mathbf{Z}_{1}=\mathcal{E}(\mathbf{X}_{1},\mathbf{A}_{1})$ and $\mathbf{Z}_{2}=\mathcal{E}(\mathbf{X}_{2},\mathbf{A}_{2})$, using the same encoder $\mathcal{E}$. Each embedding vector is then projected through a two-layer network with RELU activations to align the representations from different views while maintaining discriminative features. The contrastive loss, specifically the InfoNCE loss, is applied to align these representations while also distinguishing them from negatives within their minibatch:
@@ Appendix E Additional Results
E.1 Impact of Embedding Size on Model Performance
@@
Table 7: Impact of embedding size on link prediction performance.
@@
E.2 Performance Per Relation Type
@@
Table 8: Precision per relation type in PrimeKG++ using the Grace-Attention model.
@@
E.3 Evaluating Embedding Quality for Downstream Tasks
@@
Random
From Scratch 0.233 0
@@
LM - - 0.555 0.56
Our approach None 0.612 0.624
GGD Attention 0.646 0.656
ReDAF 0.634 0.651
None 0.621 0.601
GRACE Attention 0.636 0.611
ReDAF 0.612 0.608
None 0.633 0.625
DGI Attention 0.64 0.645
ReDAF 0.639 0.642
Table 9: Comparison of embedding methods for ML downstream task

### v2

@@ 4.4. Graph Contrastive Learning
4.5. Link Prediction in KG Embedding
KG Embedding involves an embedding function $e:E\cup R\rightarrow\mathcal{X}$, which maps entities and relations in a knowledge graph to elements within an embedding space $\mathcal{X}$. Additionally, it includes a scoring function $f:\mathcal{X}^{3}\rightarrow\mathbb{R}$ that, given the embeddings of entities and relations in a triple, computes a score indicating the likelihood or validity of the triple. In our experiment, we utilize Relational Graph Convolutional Network (RGCN) (RGCN, ) as the encoder to extract embeddings from graph-structured data that includes relational information. We then employ DistMult (dismult, ) as a scoring function to map entities and relations to vector scores. The essence of the link prediction task lies in classifying the existence of edges between entities, where positive edges are drawn from the dataset and negative edges are randomly sampled. Binary Cross Entropy (BCE) loss is employed to evaluate the effectiveness of the classification as follows:
@@ 4.5. Link Prediction in KG Embedding
$\mathcal{L}_{\text{reg}}=\lambda\left(\|\mathbf{X}\|^{2}+\|\mathbf{Z}\|^{2}\right),$
@@
To facilitate effective batch-wise training, we utilize the GraphSAINT sampling method (graphsaint, ) for graph sampling. This approach employs the Random Walks technique to sample subgraphs, ensuring the presence of existing edges within each batch for the link prediction task.
5. Experiments
5.1. Experimental Setup
@@
In our experiments, we utilize two principal datasets: PrimeKG++ and the DrugBank drug-target interaction dataset (knox2024drugbank, ). PrimeKG++ serves as our primary dataset, enriched with detailed attribute information across a variety of biological entities, making it highly suitable for comprehensive model training and evaluation. The DrugBank dataset, a curated biomedical knowledge graph, focuses specifically on drug-target protein interactions. It comprises 9,716 FDA-approved drugs and 846 protein targets, encompassing a different set of relations and nodes compared to PrimeKG++. However, the DrugBank dataset originally lacked node attributes, necessitating augmentation by incorporating detailed attribute information akin to that in PrimeKG++, thereby ensuring a comprehensive evaluation and robust performance of our model. By leveraging the enriched attribute information integrated into both datasets, we aim to thoroughly evaluate our framework’s ability to handle both broad and domain-specific biomedical knowledge graphs, enabling a rigorous assessment of its performance and generalizability.
@@
With the introduction of PrimeKG++, our augmented dataset, we conducted a comprehensive evaluation of our approach by exploring a variety of widely-used configurations. We experimented with three well-established GCL models: Graph Gaussian Diffusion (GGD), Graph Contrastive Representation Learning (GRACE), and Deep Graph Infomax (DGI). Additionally, we examined different attribute fusion methods, including Attention Fusion and Relation-guided Dual Adaptive Fusion (ReDAF), which weigh each modality differently before fusion. As a baseline, we also included a simple fusion approach (”None”) where embeddings from various modalities were combined using a mean operation without explicit weighting. To provide additional context, we compared these configurations against models trained with Random Initialization and direct Language Model (LM)-derived embeddings. Rather than focusing on identifying a single optimal configuration, our objective was to demonstrate the versatility and robustness of the proposed approach across widely-used methods. We experimented with different configurations to showcase how our framework can be applied in diverse settings. While the choice of components may depend on the specific characteristics of the dataset, our intention was to highlight the adaptability of our framework, ensuring it performs effectively under multiple configurations.
@@
The reported results are based on models with the lowest validation loss observed during training, evaluated over 100 epochs. The statistics of the dataset splits are summarized in Table 1. Our model implementations are built using PyTorch and trained on a single NVIDIA A100 GPU over 3 hours for training. This setup ensures a rigorous and reproducible evaluation framework for assessing the performance and generalizability of our proposed methods.
Table 1. Statistics of Triple Splits for PrimeKG++ and DPI Benchmark.
@@
5.2. Results and Discussion
5.2.1. PrimeKG++
@@
5.2.2. DrugBank DTI
@@
Table 2. Link prediction performance on the PrimeKG++ dataset with varying negative sampling ratios.
Initial Embedding Attribute Fusion GCL Models 1:1 1:3 1:5
@@
Direct LM-derived None - 0.993 0.975 0.982 0.934 0.972 0.902
@@
Table 3. Link prediction performance on the DrugBank DTI dataset with varying negative sampling ratios.
Initial Embedding Attribute Fusion GCL Models 1:1 1:3 1:5
@@
Direct LM-derived None - 0.994 0.957 0.988 0.884 0.982 0.822
@@
5.3. Latent Space Visualization of Embeddings
@@
Figure 3. t-SNE visualization of drug embeddings for a single protein with the highest number of interactions in the PrimeKG++ dataset. The left panel displays embeddings derived solely from the Language Model (LM), while the right panel shows embeddings generated using our proposed approach (GRACE + ReDAF). Drugs interacting with the selected protein are labeled in red, and non-interacting drugs are labeled in blue. This comparison illustrates the structural differences in the latent space resulting from the two embedding methods.
The visualization results in Figure 3 reveal notable differences between the two configurations. Embeddings generated solely with Language Models (LM) showed less distinct clustering, with considerable overlap between the two groups. This overlap suggests a limited capacity to distinguish drugs interacting with the selected protein from those that do not. In contrast, embeddings produced using our proposed method, which integrates LM with Graph Contrastive Learning (GCL), exhibited tighter clustering and more pronounced separation. This demonstrates the method’s superior ability to capture shared properties among drugs interacting with the same protein. These results underscore the robustness of our framework in generating high-quality, interpretable embeddings that accurately represent underlying biological relationships, even when applied to unseen datasets such as DrugBank.
6. Conclusion
@@ Appendix A Parameters and notations
Table 4. List of Hyperparameters
@@
Table 5. List of notations
@@
Table 6. Model summary
@@ Appendix B PrimeKG++ Creation
To obtain and integrate the metadata, in addition to PrimeKG [2], we perform the following steps:
@@
We utilize the open-source web-scraping library Selenium to fetch detailed information from authoritative sources such as Entrez Gene [21] and DrugBank [15] for each node type. This process involves the following steps: For genes/proteins, we collect both protein-coding genes, annotated with amino acid sequences, and non-coding genes, represented with nucleotide sequences. This information is fetched using the Entrez Gene API, which allows us to gather comprehensive gene data efficiently. For drugs, we source data from DrugBank, including molecular data represented as SMILES strings and antibody data identified by amino acid sequences. This comprehensive data collection ensures that we cover various aspects of drug representation. Each of these new data points is categorized into sub-types, collectively representing a node type. This detailed categorization enhances the granularity and richness of our knowledge graph.
@@ C.1. Attention Fusion
C.2. Relation-guided Dual Adaptive Fusion (ReDAF)
Given the sparse nature of PrimeKG++, we utilize the Relation-guided Dual Adaptive [40] Fusion model which produces a joint embedding projected from weighted parameters collected from individual modal training data. Additionally, the missing values of any element are consolidated with a random vector within the same vector space.
@@ Appendix D Contrastive learning settings
D.1. Deep Graph Infomax model
Deep Graph Infomax (DGI) [32] utilizes an unsupervised learning strategy for graph data by maximizing the mutual information between node representations and a global summary of the graph. The method begins with an assumption of a set of node features $X=\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N}\}$, where $N$ is the number of nodes, and each $\mathbf{x}_{i}\in\mathbb{R}^{F}$ denotes the features of node $i$. These are complemented by an adjacency matrix $A\in\mathbb{R}^{N\times N}$, which encodes the relational structure between nodes.
The core of DGI is an encoder function $E:\mathbb{R}^{N\times N}\times\mathbb{R}^{N\times F}\rightarrow\mathbb{R}^{N\times F_{0}}$ that transforms the node features and the adjacency matrix into high-level node embeddings $\{\mathbf{h}_{1},\mathbf{h}_{2},\ldots,\mathbf{h}_{N}\}$. These embeddings, or patch representations, are meant to encapsulate not only the properties of individual nodes but also their neighborhood structures.
@@ D.1. Deep Graph Infomax model
DGI employs a discriminator $D:s^{T}s$, which evaluates the mutual information between local patch representations and the global summary by assigning probability scores. These scores indicate how well the local patches (node embeddings) and the global summary correspond to each other in terms of information content.
For training, negative samples are generated through a stochastic corruption function $C:\mathbb{R}^{N\times N}\times\mathbb{R}^{N\times F}\rightarrow\mathbb{R}^{M\times M}\times\mathbb{R}^{M\times F}$, creating perturbed versions of the graph $(X_{e},A_{e})=C(X,A)$. The learning objective is to discriminate between the ”true” patch-summary pairs and those generated from corrupted inputs using a noise-contrastive estimation with a binary cross-entropy loss.
@@
D.2. Graph Group Discrimination model
We experiment with a Group-discrimination-based method called Graph Group Discrimination (GGD) [42]. Contrastive learning in this method is formulated to discriminate between groups of node embeddings, rather than individual pairs. This method leverages a binary cross-entropy loss to effectively distinguish between node samples from ‘positive’ (unaltered) and ‘negative’ (altered) graph structures.
@@ D.2. Graph Group Discrimination model
D.3. Graph Contrastive representation learning model
GRACE (Graph Contrastive Representation Learning) [43] applies stochastic augmentations to both node features and graph structure to learn robust node embeddings. For a graph with feature matrix $\mathbf{X}$ and adjacency matrix $\mathbf{A}$, two corrupted views $\mathbf{X}_{1},\mathbf{A}_{1}$ and $\mathbf{X}_{2},\mathbf{A}_{2}$ are generated by independently dropping features and edges. Node embeddings for these views are computed as $\mathbf{Z}_{1}=\mathcal{E}(\mathbf{X}_{1},\mathbf{A}_{1})$ and $\mathbf{Z}_{2}=\mathcal{E}(\mathbf{X}_{2},\mathbf{A}_{2})$, using the same encoder $\mathcal{E}$. Each embedding vector is then projected through a two-layer network with RELU activations to align the representations from different views while maintaining discriminative features. The contrastive loss, specifically the InfoNCE loss, is applied to align these representations while also distinguishing them from negatives within their minibatch:
@@ Appendix E Additional Results
E.1. Impact of Embedding Size on Model Performance
@@
Table 7. Impact of embedding size on link prediction performance.
@@
E.2. Performance Per Relation Type
@@
Table 8. Precision per relation type in PrimeKG++ using the Grace-Attention model.
@@
E.3. Evaluating Embedding Quality for Downstream Tasks
@@
Table 9. Comparison of embedding methods for ML downstream task
@@
Random From Scratch - - 0.233 0
@@
LM - - 0.555 0.560